For the piecewise function, find the values g(-7),g(2), and g(4)
g(x)= x+4, for x<2
7-x, for x>2
g(-7)=___
g(2)=___
g(4)=___
Thank you for your help

Answer by Edwin McCravy(20086)

For the piecewise function, find the values g(-7),g(2), and g(4)

g(x)= x+4, for x<2
      7-x, for x>2

There are two formulas for g(x)

1.  Use the formula g(x)=x+4 whenever x is less than 2
2.  Use the formula g(x)=7-x whenever x is greater than 2

---------------------------

g(-7)

Since -7 is less than 2 we use the first formula:

 g(x)=x+4
g(-7)=(-7)+4
g(-7)=-3

-----------------------------

g(2) 

Sorry we can't do that one because you forgot to tell us which formula
to use when x = 2, since the first one has only a <, not a <,
and the second one only has a >, not a >.  Do you forget to
underline one of the inequalities?  Since 2 is neither < 2 or > 2,
there is no way we can tell which formula to use.

------------------------------ 

g(4)

Since 4 is greater than 2 we use the second formula:

 g(x)=7-x
 g(4)=7-4
 g(4)=3

Edwin